Please use hand work and formulas as I cannot use excel and it helps me better understand. Thank you!

Assets Cash and Marketable Securities Accounts Receivable Inventories Total Current Assets Net plant and equipment TOTAL ASSETS 2017 130 4875 7995 13000 13000 26000 2016 1040 4095 5395 10530 11310 21840 Liabilities and Equity Accounts Payable Notes Payable Accruals Total Current Liabilities Long Term Bonds TOTAL DEBT Preferred Stock Common Stock Retained earnings TOTAL COMMON EQUITY TOTAL LIABILITIES AND EQUITY 2017 780 1820 1430 4030 9802 13832 520 1690 9958 11648 26000 2016 520 780 1690 2990 7540 10530 520 1690 9100 10790 21840Income Statement: 12/31/17 Net Sales Operating Costs (excludes Dep/Amortization) EBITDA Depreciation Amortization Depreciation and Amortizatiorn EBIT Less Interest EBT Taxes (40%) NET INCOME (before Preferred Dividends) Preferred Dividends NET INCOME Common Dividends Addition to Retained Earnings 2017 41600 100 6500 1300 0 1300 5200 1144 4056 622.4 2433.6 52 23816 1521 860.6 2016 37050 2461 4589 1170 1170 3419 780 2639 05.56 1583.4 52 15314 689 842.4 Refer to the statements above, compute ROE ratios for both year 2016 and 2017? Use DuPont Analysis to check for the problem of lower ROE year.

Refer to FIN300.01 Inc. What was the free cash flow in 2017 for FIN300 Inc.?
Answer #1

ROE = (Net profit / Total equity) * 100

ROE for 2016 = (1531.4 / 10790) * 100

ROE for 2016 = 14.19%

ROE for 2017 = (2381.60 / 11648) * 100

ROE for 2017 = 20.45%

DU POINT ANALYSIS FOR 2016:

Return on Equity= Net Profit Margin x Asset Turnover Ratio x Financial Leverage

ROE = (1531.40 / 37050) * (37050 / 21840) * (21840 / 10790)

ROE = .0413 * 1.69 * 2.02

ROE = 14.1%

In 2016, ROE was lower while in 2017, it increased. It came down in 2016 due to Net profit margin. Net Income is much lower in 2016 while it increased significantly in 2017.